vials pf penicillins are available in 5mL suspensions of 10mg/mL concentrations. you are asked to inject 1mg of penicillin to a patient intramuscularly. What volume of the suspension in the vial you would give to thr patient?
Answer #1

Concentration of panicillin 10mg/ ml

In 1ml suspension 10mg penicillin

In 5ml suspension 10×5 = 50mg penicillin

So 50mg penicillin in 5ml

1mg penicillin in =5/50=0.1ml solution

Answer. For 1mg penicillin we use 0.1 ml solution